Filling a tank. A water tank has an inlet pip and a drain pipe. a full tank can be emptied in 30 minutes if the drain is opened and as empty tank can be filled I 45 minutes with the inlet opened. If both pipes are accidentally opened when the tank is full, then how long will it take to empty the tank?
Answer by stanbon(48558)

Filling a tank. A water tank has an inlet pip and a drain pipe. a full tank can be emptied in 30 minutes if the drain is opened and as empty tank can be filled I 45 minutes with the inlet opened. If both pipes are accidentally opened when the tank is full, then how long will it take to empty the tank?
-----------------
Drain Pipe DATA:
Time to empty = (1/2) hr/job ; Rate = 2 job/hr
-----------
Inlet pipe DATA:
Time to fill = (3/4) hr/job ; Rate = (4/3) job/hr
------------
Working together DATA:
Rate = (4/3) - 2 = -1/3 job/hr
The tank will be emptied at (1/3) per hour.
-----------
EQUATION:
Let "x" be the number of hours to empty the tank.
x(-1/3)=-1 (-1 means 1 empty tank)
x=3 (3 hours to empty the tank)
Cheers,
Stan H.